OpenStudy (marmar10):

Help! Prove: 1-2sin^2 x/2sinxcosx = cot 2x

OpenStudy (marmar10):

\[\frac{ 1-2\sin ^{2} x}{ 2sinxcosx }= \cot 2x\]

OpenStudy (anonymous):

hi